OpenStudy (anonymous):

What is the slope intercept form of the equation 9x+3y=6

OpenStudy (anonymous):

If you don't understand any of the steps, ask!\[9x+3y=6\\3y=-9x+6\\y=-3x+2\\m=-3,y_0=2\because y=mx+y_0\]

OpenStudy (anonymous):

i dont understand what the anwser is...

OpenStudy (anonymous):

I don't specifically understand the question. A "slope intercept" doesn't mean anything, so I guess you wanted both "slope" and "intercept". Slope is generally represented as \(m\) in math; while intercept as \(y_0\).

OpenStudy (anonymous):

I get how to do the steps, but i dont get the anwser

OpenStudy (anonymous):

\[m=-3,y_0=2\]"\(\because\)" means "because".

OpenStudy (anonymous):

So \(m=-3,y_0=2\because y=mx+y_0\) is "m equals -3, y_0 equals 2, because of the form y=mx+b".